20 May 2014: Tuesday

LDAP:

  • No issues from last night’s update.
  • Work with Jim to convert mail systems to using MDB.
  • EMERGENCY update to recover from SIS feed failure.

BACKUPS:

  • Amazingly, no “misconfigured clients” after yesterday’s maintenance.  I guess that means that the list from last month is still good.  Perhaps in June, I should not run it and reset the list…
  • Clones:
    • Tape clones done at 7:43 pm yesterday (That was QUICK!).
    • Disk and CoM/IS running along with about 14TB left between them.
  • Documentation for VMware backup project.

Tags:

Leave a Reply

You must be logged in to post a comment.